Project

General

Profile

NDOS CommissionMtng 20121126

Attending: Jan, Susan, Peter, Leon, Mat, Venktesh, Vipin,

NDOS bad behavior
Test stand (lights off)

Steven Foulkes - Schedule for ioc update

Jan - Running was abysmal
- 14328 - started at 9:08 empty events, New subruns have data but only 18-30 minutes of data.
(Trigger stopping)
- 2 weeks
- Low cosmic trigger rates
- New funny readout 2-2 10, Terminator died.
- 5 dcms look like they were rebooted on Sunday at noon (5 dropped out of partitions, and were found warm.)
- Empty event 9 am and 3pm on Nov 14.

Test stand -
Leon was running DSO every 4-6 hours.
Modified the DSO script to command registers to on. (Re-enable the command registers)
Snouts in the middle of the hall way. We're seeing the hall light. Encourage people to turn off lights. Leon sees a pattern of pixels. Darkening the snouts
Need to get dcs mode running.

DaqSoftware - DCS teststand

DDF - Denis took runs with versioned data. Nearline was updated with new DDF.
- Denis want to test again
- Denis

DCS - Steven Foulkes want to update IOC to resolve sync issue.

Venktesh - Wednesday week. 8 channels of data. Synced. Today analysis.
- Flasher tests 3-6.